OpenStudy (anonymous):

It's a long one.. Given: 7x + 2 = 2 + 5x – 6 Prove: x = –3

OpenStudy (anonymous):

Given the equation 7x + 2 = 2 + 5x – 6, use the commutative property to rearrange the terms so that like terms are next to one another. This gives the equation 7x + 2 = 2 – 6 + 5x. Then use the ____________________ to group the like terms. This gives the equation 7x + 2 = (2 – 6) + 5x. Next, combine like terms to get the equation 7x + 2 = –4 + 5x. Use the subtraction property of equality to subtract 5x from both sides of the equation. This gives the equation 2x + 2 = –4. Then use the subtraction property of equality again to subtract 2 from both sides of the equation. This gives the equation 2x = –6. Finally, use the division property of equality to divide both sides of the equation by 2 to give a final solution of x = –3. Therefore, given the equation 7x + 2 = 2 + 5x – 6, x is equal to –3. Which justification was left out of the paragraph proof above?

OpenStudy (anonymous):

7x + 2 = 2 – 6 + 5x 7x - 5x = 2 -6 -2

OpenStudy (anonymous):

simplify: 2x = -6 x = -3

OpenStudy (unklerhaukus):

i think the property used is the associative of addition \[a+(b+c)=(a+b)+c\]